# Random Feature Representation Boosting

### v1

@@ 3.3 Gradient Boosting Random Features
which holds for functions $g$ with small $L_{2}^{D}(\mu)$-norm. As discussed in the context of general gradient representation boosting, a potential issue with directly minimizing $\langle g,\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)\rangle_{L^{D}_{2}(\widehat{\mu})}$ is that $g$ might learn to maximize its magnitude without following the direction of the functional gradient. To address this, we constrain the problem to ensure that $g$ maintains a unit norm in $L_{2}^{D}(\widehat{\mu})$. If we restrict $g$ to residual blocks of the form $g=\Delta f$, where $\Delta\in\mathbb{R}^{D\times p}$ is a linear map, and $f\in\mathbb{R}^{p}$ are random features, then solving the constrained $L_{2}^{D}(\widehat{\mu})$-inner product minimization problem becomes equivalent to solving a quadratically constrained least squares problem. See Appendix B for the proof.
@@
$\operatorname*{argmax}\limits_{\Delta\in\mathbb{R}^{D\times p}\textnormal{ such that }\|\Delta f\|_{L_{2}^{D}(\widehat{\mu})}\leq 1}\big{\langle}\Delta f,\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)\big{\rangle}_{L_{2}^{D}(\widehat{\mu})}$
@@
minimize $\textnormal{minimize }\quad\frac{1}{n}\sum_{i=1}^{n}\|\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)(x_{i})-\Delta f(x_{i})\|^{2},$
subject to $\textnormal{subject to }\quad\frac{1}{n}\sum_{i=1}^{n}\|\Delta f(x_{i})\|^{2}=1,$
@@
$\Delta=\frac{\sqrt{n}}{\|G\|_{Frobenius}}G^{\top}F(F^{\top}F)^{-1}$
@@
$G_{i,j}^{\textnormal{CCE}}=\big{(}W_{t-1}(p(W_{t-1}^{\top}\Phi_{t-1}(x_{i}))-e_{y_{i}})\big{)}_{j},$
where $p$ is the softmax function, and $e_{y_{i}}$ is the one-hot vector for label $y_{i}$. For full derivation, see Appendix C. We then fit $\Delta_{t}$ by fitting $f_{t+1}$ to the negative normalized functional gradient $-\frac{\sqrt{n}G}{\|G\|_{F}}$ via multiple least squares, according to Theorem B.1. In practice, we also include $\ell_{2}$ regularization.
@@
$\alpha_{t}=\operatorname*{argmin}\limits_{\alpha>0}\widehat{\mathcal{R}}(W_{t-1},\Phi_{t-1}+\alpha\Delta_{t}f_{t})$
@@
Step 3: Update the feature representation $\Phi_{t}=\Phi_{t-1}+\eta\alpha_{t}\Delta_{t}f_{t}$, and the top-level linear predictor
@@ 3.4 Theoretical Guarantees
Intuitively, this condition states that there exists a weak feature transformation in $\mathcal{G}_{t+1}$ that is negatively correlated with the functional gradient. In the sample-splitting setting, where each boosting iteration uses an independent sample of size $\widetilde{n}=\lfloor n/T\rfloor$, we derive the following regret bound for RFRBoost. This bound describes how the excess risk decays compared to an optimal predictor as $T$ and $\widetilde{n}$ vary. See Appendix D for the proof.
@@
weak feature transformations $\mathcal{G}_{t}=\{x\mapsto\Delta\tanh(B\Phi_{t-1}(x)):\lambda_{\textnormal{max}}(\Delta),\lambda_{\textnormal{max}}(B)<\lambda_{1}\}$ satisfying the $(\beta,\epsilon_{t})$-weak learning condition. Let the boosting learning rates be $\eta_{t}=ct^{-s}$ for some $s\in\big{(}\frac{\beta+1}{\beta+2},1\big{)}$ and $c>0$. Then $T$-layer RFRBoost satisfies the following risk bound for any $W^{*},\Phi^{*}$, and $a\in\big{(}0,\beta(1-s)\big{)}$ with probability at least $1-\delta$ over datasets of size $n$:
@@ 4 Numerical Experiments
In this section, we compare RFRBoost against a set of baseline models on a wide range of tabular regression and classification datasets, as well as on a challenging synthetic point cloud separation task. All our code is publicly available at github.com/nikitazozoulenko/random-feature-representation-boosting
@@ 4.1 Tabular Regression and Classification
Evaluation Procedure: We use a nested 5-fold cross-validation (CV) procedure to tune and evaluate all models, run independently for each dataset. The innermost CV is used to tune all hyperparameters, for which we use the Bayesian hyperparameter tuning library Optuna (Akiba et al., 2019). For regression, we use MSE loss, and for classification, we use cross-entropy loss. All experiments are run on a single CPU core on an institutional HPC cluster, mostly comprised of AMD EPYC 7742 nodes. We report the average test scores for each model, as well as mean training times for a single fit. The average relative rank of each model is presented in a critical difference diagram, indicating statistically significant clusters based on a Wilcoxon signed-rank test with Holm correction (Benavoli et al., 2016).
@@
Baseline Models: We compare RFRBoost, a random feature ResNet, against several strong baselines, including end-to-end (E2E) trained MLP ResNets, single-layer random feature neural networks (RFNNs), ridge regression, logistic regression, and XGBoost (Chen & Guestrin, 2016). The E2E ResNets are trained using the Adam optimizer (Kingma & Ba, 2015) with cosine learning rate annealing, ReLU activations, and batch normalization. While XGBoost is a powerful gradient boosting model, it differs fundamentally from RFRBoost. XGBoost ensembles a large number of weak decision trees (up to 1000 in our experiments) to build a strong predictor. RFRBoost, on the other hand, uses gradient representation boosting to construct a small number of residual blocks, followed by a single linear predictor. For RFRBoost and RFNNs we use SWIM random features with $\tanh$ activations, as detailed in Appendix E. In the regression setting, we evaluate three variants of RFRBoost: using a scalar, a diagonal, or a dense $\Delta$ matrix. When using a dense $\Delta$, we set the initial mapping $\Phi_{0}(x)$ to the identity; otherwise, $\Phi_{0}(x)$ is a randomly initialized dense layer. For classification, we use the gradient-greedy variant of RFRBoost.
@@
Hyperparameters: All hyperparameters are tuned with Optuna in the innermost fold, using 100 trials per outer fold, model, and dataset. For ridge and logistic regression, we tune the $\ell_{2}$ regularization. For the neural network-based models, we fix the feature dimension of each residual block to 512 and use 1 to 10 layers. For E2E networks, we tune the hidden size, learning rate, learning rate decay, number of epochs, batch size, and weight decay. For RFRBoost, we tune the $\ell_{2}$ regularization of the linear predictor and functional gradient mapping, the boosting learning rate, and the variance of the random features. For RFNNs, we tune the random feature dimension, random feature variance, and $\ell_{2}$ regularization. For XGBoost, we tune the $\ell_{1}$ and $\ell_{2}$ regularization, tree depth, boosting learning rate, and the number of weak learners. For a detailed list of hyperparameter ranges, we refer to Appendix E.
@@
Table 1: Average test RMSE and single-core CPU fit times on the OpenML regression tasks.
@@
Gradient RFRBoost 0.4086 1.749
Greedy RFRBoost $\Delta_{dense}$ 0.4094 2.888
Greedy RFRBoost $\Delta_{diag}$ 0.4152 0.918
Greedy RFRBoost $\Delta_{scalar}$ 0.4333 0.638
XGBoost 0.3954 1.873
E2E MLP ResNet 0.4167 38.909
RFNN 0.4340 0.044
Ridge Regression 0.5395 0.001
Figure 2: Critical difference diagram based on relative rank of test RMSE. Bars indicate no significant difference ($\alpha=0.10$).
Results: Regression and classification results can be found in Tables 1 to 2 and Figures 2 to 3, respectively. We find that RFRBoost ranks significantly better than all other baseline models. For regression tasks, the gradient-greedy version of RFRBoost outperforms the exact-greedy variant, contrary to the observations of Suggala et al. (2020) for SGD-trained gradient representation boosting. This difference likely arises because the SGD-based approach does not incorporate the $L_{2}^{D}(\mu)$-norm constraint during training, which is crucial for preserving the functional direction of the residual block. The test scores follow the ordering $\Delta_{\textnormal{dense}}>\Delta_{\textnormal{diag}}>\Delta_{\textnormal{scalar}}$, demonstrating that RFRBoost is more expressive when mapping random feature layers to the functional gradient, rather than simply stacking random feature layers. While XGBoost achieves a slightly lower RMSE than RFRBoost, it performs worse in terms of average rank. Moreover, RFRBoost significantly outperforms both RFNNs and E2E MLP ResNets, while being an order of magnitude faster to train than the latter. Although the reported training times are CPU-based, our implementation suggests both methods would benefit similarly from GPU acceleration, making the presented times representative.
@@
RFRBoost 0.854 2.006
XGBoost 0.852 3.695
E2E MLP ResNet 0.850 30.518
RFNN 0.845 1.022
Logistic Regression 0.821 0.134
Figure 3: Critical difference diagram based on relative rank of test accuracy. Bars indicate no significant difference ($\alpha=0.10$).
4.2 Point cloud separation
We evaluate RFRBoost on a challenging synthetic dataset originating from the neural ODE literature (Sander et al., 2021). The task is to linearly separate (i.e. classify) a set of concentric circles while restricting the ResNet hidden size to 2, see Figure 4. We compare RFRBoost to E2E MLP ResNets, while also presenting classification results for logistic regression and RFNNs in Table 3. All models were trained with cross entropy loss, using a 5-fold CV grid search for hyperparameter tuning. For E2E MLP ResNets, the learning rate was tuned, while for RFRBoost, only the $\ell_{2}$ regularization of the gradient mapping was tuned. The hidden size, residual block feature dimension, and activation function was fixed at $2$, $512$, and $\tanh$ respectively, for all models. See Appendix E for more details.
@@ 4.2 Point cloud separation
RFNN 0.878 0.027
E2E MLP ResNet 0.519 0.121
Logistic Regression 0.327 0.019
@@
Notably, the RFNN, which uses SWIM random features to map the initial 2-dimensional input to a higher-dimensional space before applying a linear classifier, fails to classify all points correctly. RFRBoost, in contrast, achieves near-perfect linear separation by first mapping the random features to the functional gradient of the network representation, before applying a 2-dimensional linear classifier. This result is somewhat surprising because RFRBoost does not rely on explicit Fourier features or a learned radial basis, demonstrating the power of gradient representation boosting. It further illustrates how RFRBoost can solve problems that are intractable for E2E-trained networks and RFNNs. Figure 4 shows that the E2E MLP ResNet correctly separates only three out of the nine concentric rings, similar to the failure to converge observed by Sander et al. (2021).
@@ 5 Conclusion
This paper introduced RFRBoost, a novel method for constructing deep residual random feature neural networks (RFNNs) using boosting theory. RFRBoost addresses the limitations of single-layer RFNNs by using random features to learn a feature representation that approximates the functional gradient of the network, thereby enhancing performance while retaining the computational benefits of convex optimization for RFNNs. We derived closed-form solutions for greedy layer-wise boosting with MSE loss, and presented a general framework for other loss functions based on quadratically constrained least squares. Through numerical experiments on tabular data, we demonstrated that RFRBoost significantly outperforms both traditional RFNNs and end-to-end trained MLP ResNets, while offering substantial computational advantages and theoretical generalization guarantees. RFRBoost represents a significant step towards building powerful, stable, efficient, and theoretically sound deep networks using random features. Future work will focus on extending RFRBoost to other domains such as time series data, exploring different types of random features and momentum strategies, implementing efficient GPU acceleration, and scaling to large datasets.

### v2

@@ 3.3 Gradient Boosting Random Features
which holds for functions $g$ with small $L_{2}^{D}(\mu)$-norm. As discussed in the context of general gradient representation boosting, a potential issue with directly minimizing $\langle g,\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)\rangle_{L^{D}_{2}(\widehat{\mu})}$ is that $g$ might learn to maximize its magnitude without following the direction of the functional gradient. To address this, we constrain the problem to ensure that $g$ maintains a unit norm in $L_{2}^{D}(\widehat{\mu})$. If we restrict $g$ to residual blocks of the form $g=Af$, where $A\in\mathbb{R}^{D\times p}$ is a linear map, and $f\in\mathbb{R}^{p}$ are random features, then solving the constrained $L_{2}^{D}(\widehat{\mu})$-inner product minimization problem becomes equivalent to solving a quadratically constrained least squares problem. See Appendix B for the proof.
@@
$\operatorname*{argmax}\limits_{A\in\mathbb{R}^{D\times p}\textnormal{ such that }\|Af\|_{L_{2}^{D}(\widehat{\mu})}\leq 1}\big{\langle}Af,\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)\big{\rangle}_{L_{2}^{D}(\widehat{\mu})}$
@@
minimize $\textnormal{minimize }\quad\frac{1}{n}\sum_{i=1}^{n}\|\nabla_{2}\widehat{\mathcal{R}}(W,\Phi)(x_{i})-Af(x_{i})\|^{2},$
subject to $\textnormal{subject to }\quad\frac{1}{n}\sum_{i=1}^{n}\|Af(x_{i})\|^{2}=1,$
@@
$A=\frac{\sqrt{n}}{\|G\|_{Frobenius}}G^{\top}F(F^{\top}F)^{-1}$
@@
$G_{i,j}^{\textnormal{CCE}}=\big{(}W_{t-1}(s(W_{t-1}^{\top}\Phi_{t-1}(x_{i}))-e_{y_{i}})\big{)}_{j},$
where $s$ is the softmax function, and $e_{y_{i}}$ is the one-hot vector for label $y_{i}$. For full derivation, see Appendix C. We then fit $A_{t}$ by fitting $f_{t+1}$ to the negative normalized functional gradient $-\frac{\sqrt{n}G}{\|G\|_{F}}$ via multiple least squares, according to Theorem B.1. In practice, we also include $\ell_{2}$ regularization.
@@
$\alpha_{t}=\operatorname*{argmin}\limits_{\alpha>0}\widehat{\mathcal{R}}(W_{t-1},\Phi_{t-1}+\alpha A_{t}f_{t})$
@@
Step 3: Update the feature representation $\Phi_{t}=\Phi_{t-1}+\eta\alpha_{t}A_{t}f_{t}$, and the top-level linear predictor
@@ 3.4 Theoretical Guarantees
Intuitively, this condition states that there exists a weak feature transformation in $\mathcal{G}_{t+1}$ that is negatively correlated with the functional gradient. In the sample-splitting setting, where each boosting iteration uses an independent sample of size $\widetilde{n}=\lfloor n/T\rfloor$, we derive the following regret bound for RFRBoost. This bound describes how the excess risk decays compared to an optimal predictor as $T$ and $\widetilde{n}$ vary.
@@
weak feature transformations $\mathcal{G}_{t}=\{x\mapsto A\tanh(B\Phi_{t-1}(x)):\lambda_{\textnormal{max}}(A),\lambda_{\textnormal{max}}(B)<\lambda_{1}\}$ satisfying the $(\beta,\epsilon_{t})$-weak learning condition. Let the boosting learning rates be $\eta_{t}=ct^{-s}$ for some $s\in\big{(}\frac{\beta+1}{\beta+2},1\big{)}$ and $c>0$. Then $T$-layer RFRBoost satisfies the following risk bound for any $W^{*},\Phi^{*}$, and $a\in\big{(}0,\beta(1-s)\big{)}$ with probability at least $1-\delta$ over datasets of size $n$:
@@
Proof.
See Appendix D. ∎
For similar results in the literature, see for instance Suggala et al. (2020), Corollary 4.3.
3.5 Time Complexity
The serial time complexity of RFRBoost with MSE loss, assuming tabular data and dense random features, is $O(T[N(D^{2}+Dd+p^{2}+pD)+D^{3}+dD^{2}+p^{3}+Dp^{2}])$, as derived from Algorithm 2. Here $N$ is the dataset size, $D$ is the dimension of the neural network representation, $d$ is the output dimension of the regression task, $p$ is the number of random features, and $T$ is the number of layers (boosting rounds). The computation is dominated by matrix operations, which are well-suited for GPU acceleration. The classification case follows similarly.
@@ 4 Numerical Experiments
In this section, we compare RFRBoost against a set of baseline models on a wide range of tabular regression and classification datasets, as well as on a challenging synthetic point cloud separation task. All our code is publicly available at https://github.com/nikitazozoulenko/random-feature-representation-boosting.
@@ 4.1 Tabular Regression and Classification
Evaluation Procedure: We use a nested 5-fold cross-validation (CV) procedure to tune and evaluate all models, run independently for each dataset. The innermost CV is used to tune all hyperparameters, for which we use the Bayesian hyperparameter tuning library Optuna (Akiba et al., 2019). For regression, we use MSE loss, and for classification, we use cross-entropy loss. All experiments are run on a single CPU core on an institutional HPC cluster, mostly comprised of AMD EPYC 7742 nodes. We report the average test scores for each model, as well as mean training times for a single fit. The average relative rank of each model is presented in a critical difference diagram, indicating statistically significant clusters based on a Wilcoxon signed-rank test with Holm correction. This is a well-established methodology for comparing multiple algorithms across multiple datasets (Demšar, 2006; García & Herrera, 2008; Benavoli et al., 2016).
@@
Baseline Models: We compare RFRBoost, a random feature ResNet, against several strong baselines, including end-to-end (E2E) trained MLP ResNets, single-layer random feature neural networks (RFNNs), ridge regression, logistic regression, and XGBoost (Chen & Guestrin, 2016). The E2E ResNets are trained using the Adam optimizer (Kingma & Ba, 2015) with cosine learning rate annealing, ReLU activations, and batch normalization. While XGBoost is a powerful gradient boosting model, it differs fundamentally from RFRBoost. XGBoost ensembles a large number of weak decision trees (up to 1000 in our experiments) to build a strong predictor. RFRBoost, on the other hand, uses gradient representation boosting to construct a small number of residual blocks, followed by a single linear predictor. For RFRBoost and RFNNs we use SWIM random features with $\tanh$ activations, as detailed in Appendix E. In the regression setting, we evaluate three variants of RFRBoost: using a scalar, a diagonal, or a dense $A$ matrix. When using a dense $A$, we set the initial mapping $\Phi_{0}(x)$ to the identity; otherwise, $\Phi_{0}(x)$ is a randomly initialized dense layer. For classification, we use the gradient-greedy variant of RFRBoost with LBFG-S as the convex solver. We use the official implementation of XGBoost, and implement all other baseline models in PyTorch (Paszke et al., 2019).
@@
Hyperparameters: All hyperparameters are tuned with Optuna in the innermost fold, using 100 trials per outer fold, model, and dataset. For ridge and logistic regression, we tune the $\ell_{2}$ regularization. For the neural network-based models, we fix the feature dimension of each residual block to 512 and use 1 to 10 layers. For E2E networks, we tune the hidden size, learning rate, learning rate decay, number of epochs, batch size, and weight decay. For RFRBoost, we tune the $\ell_{2}$ regularization of the linear predictor and functional gradient mapping, the boosting learning rate, and the variance of the random features. For RFNNs, we tune the random feature dimension, random feature variance, and $\ell_{2}$ regularization. For XGBoost, we tune the $\ell_{1}$ and $\ell_{2}$ regularization, tree depth, boosting learning rate, and the number of weak learners. For a detailed list of hyperparameter ranges, along with an ablation study comparing SWIM random features to i.i.d. Gaussian random features, we refer the reader to Appendix E.
Results: Summary results for regression and classification are presented in Tables 1 to 2 and Figures 2 to 3, respectively. Full dataset-wise results are reported in Appendix E. We find that RFRBoost ranks higher than all other baseline models. For regression tasks, the gradient-greedy version of RFRBoost outperforms the exact-greedy variant, contrary to the observations of Suggala et al. (2020) for SGD-trained gradient representation boosting. This difference likely arises because the SGD-based approach does not incorporate the $L_{2}^{D}(\mu)$-norm constraint during training, which is crucial for preserving the functional direction of the residual block. The test scores follow the ordering $A_{\textnormal{dense}}>A_{\textnormal{diag}}>A_{\textnormal{scalar}}$, demonstrating that RFRBoost is more expressive when mapping random feature layers to the functional gradient, rather than simply stacking random feature layers. While XGBoost achieves a slightly lower RMSE than RFRBoost, it performs worse in terms of average rank. Moreover, RFRBoost significantly outperforms both RFNNs and E2E MLP ResNets, while being an order of magnitude faster to train than the latter. Although the reported training times are CPU-based, our implementation suggests both methods would benefit similarly from GPU acceleration, making the presented times representative.
@@
Table 1: Average test RMSE and single-core CPU fit times on the OpenML regression datasets.
@@
Gradient RFRBoost 0.408 1.688
Greedy RFRBoost $A_{dense}$ 0.408 2.734
Greedy RFRBoost $A_{diag}$ 0.415 1.631
Greedy RFRBoost $A_{scalar}$ 0.434 1.024
XGBoost 0.394 1.958
E2E MLP ResNet 0.412 19.309
RFNN 0.434 0.053
Ridge Regression 0.540 0.001
Figure 2: Critical difference diagram based on pairwise relative rank of test RMSE. Bars indicate no significant difference ($\alpha=0.05$). The average rank is displayed for each model.
@@
RFRBoost 0.853 2.519
XGBoost 0.853 3.859
E2E MLP ResNet 0.851 20.881
RFNN 0.845 1.189
Logistic Regression 0.821 0.165
Figure 3: Critical difference diagram based on pairwise relative rank of test accuracy. Bars indicate no significant difference ($\alpha=0.05$). The average rank is displayed for each model.
@@
RFNN 0.887 0.037
E2E MLP ResNet 0.732 0.144
Logistic Regression 0.334 0.023
@@
4.2 Point Cloud Separation
We evaluate RFRBoost on a challenging synthetic dataset originating from the neural ODE literature (Sander et al., 2021). The dataset consists of 10,000 points sampled from concentric circles, and the task is to linearly separate (i.e. classify) the concentric circles while restricting the ResNet hidden size to 2, see Figure 4. We compare RFRBoost to E2E MLP ResNets, while also presenting classification results for logistic regression and RFNNs in Table 3. All models were trained with cross entropy loss, using a 5-fold CV grid search for hyperparameter tuning. For E2E MLP ResNets, the learning rate was tuned, while for the other models, only the $\ell_{2}$ regularization of the classification head was tuned. The hidden size, residual block feature dimension, and activation function was fixed at $2$, $512$, and $\tanh$, respectively, for all models. See Section E.3 for more details.
Notably, the RFNN, which uses SWIM random features to map the initial 2-dimensional input to a higher-dimensional space before applying a linear classifier, fails to classify all points correctly. RFRBoost, in contrast, achieves near-perfect linear separation by first mapping the random features to the functional gradient of the network representation, before applying a 2-dimensional linear classifier. This result is somewhat surprising because RFRBoost does not rely on explicit Fourier features or a learnt radial basis, demonstrating the power of gradient representation boosting. It further illustrates how RFRBoost can solve problems that are intractable for E2E-trained networks and RFNNs. Figure 4 shows that the E2E MLP ResNet correctly separates only two of the nine concentric rings, similar to the failure to converge observed by Sander et al. (2021).
4.3 Experiments on Larger-scale Datasets
To complement our experiments on the OpenML benchmark suite, we conducted additional full-scale evaluations on four larger datasets (two with $\sim$100k samples, two with $\sim$500k samples) to assess performance and scalability in larger data regimes. Full experimental details, including dataset splits, hyperparameter grids, and plots of training time and predictive performance versus training set size, are provided in Section E.5. We find that RFRBoost significantly outperforms traditional single-layer RFNNs across all tested datasets and training sizes, particularly as dataset size increases, highlighting its effectiveness in leveraging depth for random feature models. While RFRBoost performs strongly against E2E trained networks in medium-sized data regimes, our findings indicate that E2E networks and XGBoost eventually outperform RFRBoost as dataset size increases on 3 out of 4, and 2 out of 4 datasets, respectively. We hypothesize that using RFRBoost as an initialization strategy, followed by end-to-end fine-tuning, could be a promising direction to further enhance the performance of deep networks. We leave this to future work.
@@ 5 Conclusion
This paper introduced RFRBoost, a novel method for constructing deep residual random feature neural networks (RFNNs) using boosting theory. RFRBoost addresses the limitations of single-layer RFNNs by using random features to learn a feature representation that approximates the functional gradient of the network, thereby enhancing performance while retaining the computational benefits of convex optimization for RFNNs. We derived closed-form solutions for greedy layer-wise boosting with MSE loss, and presented a general framework for arbitrary loss functions based on solving a quadratically constrained least square problem. Through extensive numerical experiments on tabular datasets for both regression and classification, we have demonstrated that RFRBoost significantly outperforms traditional RFNNs and end-to-end trained MLP ResNets in the small- to medium-scale regime where RFNNs are typically applied, while offering substantial computational advantages, and theoretical guarantees stemming from boosting theory. RFRBoost represents a significant step towards building powerful, stable, efficient, and theoretically sound deep networks using untrained random features. Future work will focus on extending RFRBoost to other domains such as time series or image data, exploring different types of random features and momentum strategies, implementing more efficient GPU acceleration, scaling to large datasets, and using RFRBoost as an initialization strategy for large-scale end-to-end training.